# COMPACT: COMPositional Atomic-to-Complex Visual Capability Tuning

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) excel at simple vision-language tasks but struggle when faced with complex tasks that require multiple capabilities, such as simultaneously recognizing objects, counting them, and understanding their spatial relationships. This might be partially the result of the fact that Visual Instruction Tuning (VIT), a critical training step for MLLMs, has traditionally focused on scaling data volume, but not the compositional complexity of training examples. We propose COMPACT (COMPositional Atomic-to-complex Visual Capability Tuning), which generates a training dataset explicitly controlling for the compositional complexity of the training examples. The data from COMPACT allows MLLMs to train on combinations of atomic capabilities to learn complex capabilities more efficiently. Across all benchmarks, COMPACT achieves comparable performance to the LLaVA-665K VIT while using less than 10% of its data budget, and even outperforms it on several, especially those involving complex multi-capability tasks. For example, COMPACT achieves substantial 83.3% improvement on MMStar and 94.0% improvement on MM-Vet compared to the full-scale VIT on particularly complex questions that require four or more atomic capabilities. COMPACT offers a scalable, data-efficient, visual compositional tuning recipe to improve on complex visual-language tasks.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have shown impressive progress in handling a wide range of vision-language tasks [2, 1, 19]. Yet, from early diagnostic works [30, 32] to recent state-of-the-art models like LLaVA [22, 23], Cambrian [37], and Eagle [34, 21], the compositionality challenge remains persistent. Consider the following question: “Are there more blue squares or red circles on the image?” A model that is capable of recognizing shapes, colors and counting objects should be able to answer it correctly. Despite years of progress, state-of-the-art models still fail on such compositional questions, even though they can answer simpler ones correctly (e.g. “What color is the square?”). This has been a long‑standing issue and such failures suggest that current models do not systematically generalize to tasks with higher compositional complexity.
To address this, recent efforts have primarily scaled up the amount of training data used for Visual Instruction Tuning (VIT) [37, 24, 25, 34, 21], an essential but data- and compute-heavy step for MLLM training. However, such datasets (e.g. LLaVA-665K [24]) are dominated by simple queries that require only one capability, lacking sufficient compositional complexity (Fig. 1). Even with large-scale instruction tuning, recent studies show that models still struggle with integrating capabilities and generalizing to complex visual tasks due to limitations in the compositional complexity of their training data [29, 40]. Thus, we ask: Can we bridge the performance gap between different complexity regimes without simply increasing the data volume?
Figure 1: Compositional Complexity Comparison. Comparison between visual instruction tuning data (LLaVA-665K [24] VIT) and our compositional tuning data (COMPACT). The VIT data is dominated by simple queries ($k=1$), while our COMPACT data is balanced across compositional complexity levels ($k=1,2,3$).
Instead of treating compositionality as a byproduct of scale, we encourage compositional capabilities in MLLMs with hierarchically structured compositional training data. In this work, we introduce COMPACT (COMPositional Atomic-to-complex Visual Capability Tuning), a data recipe that scales capabilities of MLLMs from atomic ($k=1$) to composite ($k>1$) complexity levels. We define a set of 10 atomic capabilities and then combine them to generate a compositional training dataset that can promote model’s internalization of the compositional structures of complex tasks in a compute-efficient manner. We summarize our key contributions:
  • We introduce COMPACT, a visual compositional tuning data recipe that builds complex capabilities from simple atomic capabilities. By systematically combining 10 atomic capabilities to control the complexity of training samples, COMPACT addresses a key limitation of conventional VIT methods that rely on incidental capability composition through data scaling.
  • We develop a structured data recipe that enforces a balanced distribution across different levels of compositional complexity ($k=1,2,3$) to cover a wider range of task regimes. This approach flattens the complexity cliff in conventional VIT datasets [24], where 90.1% of the questions require two or fewer capabilities.
  • We demonstrate the impressive effectiveness of COMPACT. With only 10% of the size of the VIT dataset (5% of the LLaVA-665K [24] VIT data augmented with 32K samples of our compositional tuning data), training with COMPACT matches the performance of full-scale VIT (100.18% relative score). Further, it demonstrates exceptional generalization to higher-complexity tasks, improving the score from 35.3 to 64.7 on MMStar [6] and from 32.5 to 57.5 on MMVet [43] for $k=4$ tasks.
Figure 2: COMPACT’s Data Generation Pipeline. (Left): We sample $k\in\{1,2,3\}$ atomic capabilities such as color, object recognition, and spatial relationship. (Center): We generate questions that integrate all $k$ sampled capabilities. (Right): We verify the quality of generated conversations and combine them with instruction tuning data to maintain instruction following capability. This structured data recipe explicitly models atomic-to-complex learning procedure, in contrast to standard LLaVA-665K [24] VIT that promotes learning from simple queries.
2 Related Work
Visual Instruction Tuning.  Instruction following is an essential capability in language models [38, 46]. Misalignment between a model’s response and the format requested by a question can hinder the precise evaluation of its performance and capabilities [11, 12, 31, 3]. In order to adapt MLLMs to respond appropriately to diverse question formats (e.g., multiple-choice, short- and long-response questions), Visual instruction tuning [22, 24] has been proposed. VIT involves training a model on a fixed set of instruction patterns that can be repeated during inference. Although VIT has shown performance improvements in general multimodal capabilities [14], recent work [10] has shown that optimizing for response formatting potentially limits the quality of language model responses.
While VIT [24] focuses on learning simple capabilities via instruction following, our data recipe explicitly models compositional capabilities in the training data. Our approach directly addresses the lack of exposure to compositional questions during training, enabling models to improve on tasks that are more complex in capability space. COMPACT leverages the learning potential of both compositional tuning and instruction tuning data to create a more optimal data recipe than conventional VIT.
Compositionality in LLMs and MLLMs.  Semantically, compositionality is the claim that the meaning of a complex statement is a result of the combination of its constituents. [8]. In the context of visual capabilities of MLLMs, compositional capability refers to a model’s ability to perform complex tasks by combining multiple capabilities [13], where each capability is related to understanding basic visual concepts such as objects, attributes, relationships. Recent work has shown that compositional capability can be trained in LLMs [42, 45], but generalizations to the realms of MLLMs have been largely incomplete. Some studies highlight that while MLLMs do show signs of compositional capability [28], they struggle when constituting components and their combined patterns are not strongly learned or missing during training [4]. Furthermore, previous works have focused on limited domains such as geometry [5], visual recognition, and language [7], or employed a relaxed definition of compositionality as a sequential array of tasks [20] rather than integrating them.
Studies show that general visual capability requires strong compositional ability [44]. In order to train MLLMs to learn complex capabilities, it is necessary to explicitly model compositionality in the training data. Our approach takes advantage of these findings to create a data recipe for training complex capabilities across visual domains.
Data Efficiency in MLLMs.  VIT is a data and compute-heavy step in training [41]. Studies have found that the performance of MLLMs can be reproduced with less data and better techniques, suggesting that the amount of data needed for VIT can be reduced. For example, recent works have developed effective VIT data recipes by leveraging data selection methods and curating higher-quality training datasets [16, 26]. ICONS [39] shows that models can achieve near-perfect performance across a suite of MLLM benchmarks with a fraction of the original VIT dataset. On the other hand, some studies proposed an alternative approach of scaling up to improve visual capabilities even further [18].
However, these approaches treat compositionality as a byproduct of scale rather than as a learnable capability. Our COMPACT formalizes atomic capabilities and systematically incorporates their combinations into the training dataset to efficiently address the limitations in generalization to complex compositional tasks. By redistributing the compositional complexity of the training data, we scale the model’s exposure to complex tasks without scaling the data.
3 Method
We propose COMPACT (Fig. 2), a data recipe that scales capabilities of MLLMs from atomic ($k=1$) to composite ($k>1$) complexity levels. We first introduce the concept of atomic visual capabilities (§3.1), which serve as the foundational building blocks of complex visual tasks. We then detail COMPACT’s four-step data recipe (§3.2) for generating high-quality compositional training data that allows models to integrate multiple visual capabilities.
3.1 Atomic Visual Capabilities
Atomic capabilities are foundational skills that can be combined to solve complex tasks. For example, a model needs to acquire object recognition, color attribution, and spatial relationship understanding capabilities to identify how objects of different colors are spatially oriented. For each task $T$, we identify a set of atomic visual capabilities $\{c_{1},\dots c_{k}\}$ required to solve this task. We define the number of atomic capabilities required to solve the task $T$ as its compositional complexity $k$. Our goal in COMPACT is to increase the average compositional complexity of the dataset and balance its distribution among the samples so that each question explicitly requires the model to combine multiple atomic capabilities during training.
@@ 3.1 Atomic Visual Capabilities
Table 1: Taxonomy of Atomic Capabilities. We identify 10 atomic capabilities and categorize them into three groups: Attribution, Recognition, and Relation. Atomic capabilities serve as building blocks for compositional instruction tuning. For each capability, we provide the definition and a question example that requires the capability to answer.
@@
We build a taxonomy of atomic capabilities from the existing literature on MLLMs and their general visual reasoning skills [15, 40]. Extremely low-frequency and non-perceptual capabilities (e.g. cultural knowledge, historical context, and math) are removed, resulting in 10 fine-grained atomic capabilities (Tab. 1) that focus on visual understanding. We categorize these atomic capabilities into three major categories: 1) Attribution: identifying visual properties of objects (e.g., color and shape). 2) Recognition: detecting and interpreting visual entities, including objects, actions, text, spatial recognition, and counts. 3) Relation: capturing how entities interact or relate to one another from an abstract or a physical perspective, either through spatial relationship, object interaction, or scene understanding.
@@ 3.2 Visual Compositional Tuning Data Recipe
In our proposed approach COMPACT, we generate multi-capability questions $\mathcal{D}_{\text{comp}}$ by prompting vision-language models to create questions that require natural [^1] We use the term “natural” to denote combination of visual capabilities that correspond to their co-occurrence patterns in real-world settings, wherein multiple capabilities are integrated in a way that is contextually and semantically meaningful. integration of exactly $k$ atomic visual capabilities. This process involves four key steps: (1) randomly selecting $k$ capabilities from our predefined set of atomic visual capabilities (Capability Sampling), (2) prompting Gemini-2.0-Flash [36] to generate questions that naturally integrate all $k$ selected capabilities (Conversation Generation), (3) validating the capability requirement and the quality of each question through an independent verification step (Quality Verification), and (4) combining our generated compositional tuning data with a small portion of the LLaVA-665K [24] VIT data to assemble the final dataset (Dataset Assembly).
@@
Step 1: Capability Sampling.  We start by taking a random sample of images from LLaVA-665K [24]. For each image, we repeatedly sample $k\in\{1,2,3\}$ capabilities from our predefined pool of 10 atomic visual capabilities. In order to include diverse combinations of atomic capabilities in our training dataset, we do the following in each round of capability sampling: (a) prioritize the capabilities that have not been selected for that image, and (b) drop duplicate combinations of capabilities for the same image. These efforts ensure that our training examples efficiently capture the visual information in the images.
@@
Step 2: Conversation Generation.  For each capability combination that is sampled, we prompt Gemini-2.0-Flash [36] to generate a conversational question-answer pair that integrates all capabilities in the combination, as well as a score between 0 and 100 that represents its confidence in the quality of the conversation. Our carefully designed prompt (see Appendix §B) enforces several key constraints: (a) questions must require the use of visual information from the image and cannot be answered from its text alone, (b) answers must be concise, (c) questions must integrate exactly the specified capabilities naturally (without using conjunctions to simply conjoin single-capability questions), and (d) questions must reference objects and features actually present in the image. The purpose of these constraints is to produce vision-centric conversations that are unambiguous and natural. The format of the generated output adheres to a JSON template that explicitly tags the required capabilities for each question.
@@
Step 3: Quality Verification.  We include a verification process with Gemini-2.0-Flash [36] to ensure the quality and diversity of the training dataset. We filter out questions with uninformative answers (e.g., “unknown”, “not visible”) or those with confidence scores below 70%. We discard questions that share more than 60% of the words with those previously accepted. Additionally, we reject visually ungrounded questions whose answers can be inferred from the question alone.
Then, we perform capability verification by prompting Gemini-2.0-Flash [36] to analyze whether each question indeed requires exactly the $k$ specified capabilities. Questions that require unspecified capabilities or do not utilize all $k$ capabilities are rejected. The generation and verification processes in steps 2 and 3 repeat iteratively until we collect 2-3 high-quality conversations per $k$ for each image or reach a maximum of 10 verification attempts. Only images with at least two verified questions are included in the final dataset. Step 4: Dataset Assembly.  The final training dataset combines two components: (1) a random 5% subset of the LLaVA-665K [24] VIT dataset, and (2) our COMPACT-generated compositional tuning data. The compositional tuning data consists of compositional multiturn conversations and their corresponding images randomly sampled from the VIT dataset. This mixture of instruction tuning and compositional tuning data serves a dual purpose. First, the VIT subset maintains the model’s ability to handle diverse response formats and instructions required by modern MLLM benchmarks (e.g., multiple-choice questions [9], open-ended answers [24]). Second, our compositional data trains the model’s capability to reason about multiple visual aspects within a single complex question. In this way, we delegate the instruction following capability training to the original VIT data and allow our compositional tuning data to focus on developing the model’s compositional capabilities.
COMPACT preserves the contents of the images sampled from LLaVA-665K [24] when generating new multi-turn conversations. This enables us to fairly compare COMPACT and existing methods in their ability to extract rich and structured information from the controlled set of images. We further adjust the ratio of the VIT subset to our compositional tuning data and study how it affects performance (§4.4). Our findings show the optimal balance between the preservation of instruction following capability and the training of compositional capabilities.
@@
Recipe # Data InfoVQA [27] SeedBench2Plus [17] MME [9] TextVQA [35] MM-Vet [43] CV-Bench [37] MMStar [6] LLaVA-W [24] Rel. (%)
LLaVA-665K [24] 665K 20.80 41.72 1478.48 46.99 29.22 60.92 35.11 68.50 100.00
@@
ICONS [39] 65K 21.0 42.03 1402.75 43.12 31.23 55.96 35.96 61.8 97.47
@@
Table 2: Baseline Comparisons. Performance comparison of COMPACT with baselines. With only 5% of the LLaVA-665K [24] VIT data and 32K of our compositional tuning data (65K total), COMPACT outperforms the random subset of the VIT data (Random), gradient-based approach selected subset of the VIT data (ICONS [39]), and even the full VIT data on diverse multimodal benchmarks. The best and second best results for each benchmark are shown in bold and underlined, respectively. COMPACT integrates atomic capabilities into tasks of higher compositional complexity, enabling models to generalize and handle complex tasks without explicit decomposition.

### v2

@@ Abstract
Visual instruction tuning (VIT) datasets are constructed from randomly sampled image-question pairs, without regard to the informativeness of each pair. Recent dataset selection methods have shown that a small fraction of such datasets enriched with informative samples can lead to efficient finetuning of Multimodal Large Language Models. In this work, we explore the impact of sample complexity on informative data curation and introduce COMPACT (COMPositional Atomic-to-complex Visual Capability Tuning), a VIT data recipe that scales training sample complexity by combining multiple atomic visual capabilities in a single training example. Concretely, we synthesize rich and informative text questions for each image, allowing us to significantly reduce the number of training examples required for effective visual instruction tuning. COMPACT demonstrates superior data efficiency compared to existing data reduction methods. When applied to the LLaVA-665K VIT dataset, COMPACT reduces the data budget by 90% while still achieving 100.2% of the full VIT performance (compared to only 97.5% by the state-of-the-art method) across eight multimodal benchmarks. Further, training on the COMPACT data outperforms training on the full-scale data on particularly complex benchmarks such as MM-Vet (+8.6%) and MMStar (+2.9%). COMPACT offers a scalable and efficient synthetic data generation recipe to improve on visual language tasks.
@@ 1 Introduction
Visual instruction tuning (VIT) data for Multimodal Large Language Models (MLLMs) have continuously scaled over time. Cambrian-10M (cvbench) is 15 times larger than LLaVA-665K (liu2024improved), and Eagle 2 (li2025eagle) instruction tuning data is 2.6 times larger than Cambrian-10M. As a result, state-of-the-art MLLMs like LLaVA (liu2023visual; liu2024improved), Cambrian (cvbench), and Eagle (shi2024eagle; li2025eagle) have shown impressive progress in a wide range of vision-language tasks (bai2023qwen; alayrac2022flamingo; li2022blip). The prevailing axiom of quantity over quality (li2024llava) in multimodal training shifted the attention of subsequent works away from the fundamental question: Can we develop more effective data curation methods beyond blind scaling?
Visual reasoning often relies on combining multiple fundamental visual capabilities (ke2025explain; wu2024conceptmix; zerroug2022benchmark). Consider the question, “What color is the object on the left side of the car?”. A model needs to see the car (object recognition), find what is on the left side (spatial relationship), and identify its color (color attribution) to answer the question. However, VIT datasets typically focus on individual or limited combinations of visual capabilities (e.g. “What color is the car?”), ignoring the crucial relationship between these capabilities and how they might be combined to solve complex tasks. The resulting visual reasoning questions are often simplistic and refer to a limited region in the image, under-utilizing the rich visual information within. We therefore ask whether scaling the complexity of each sample can lead to more informative datasets.
We first define a set of fundamental visual capabilities essential for visual reasoning, called atomic capabilities (Tab. 1), by analyzing the LLaVA-665K VIT dataset (see §3.1 and §A.4 for more details on the curation process). Using these as building blocks, we define a complexity measure, called $k$-value, as the number of atomic capabilities required to answer a question. This measure allows us to quantify the complexity of each sample and identify combinations of capabilities that improve information density. For example, instead of the $k=2$ question, “What color is the car?” (object recognition and color attribution), we can ask a $k=3$ question, “What color is the object on the left side of the car?” (object recognition, color attribution, and spatial understanding). Analysis of existing VIT datasets reveals an over-representation of simpler ($k\leq 2$) questions that under-exploit the visual information in images.
We conduct an exploratory experiment where we take a small amount of VIT data and increase the $k$-value for a subset of examples. We first identify the atomic visual capabilities required to answer each question. We then regenerate the questions with Gemini-2.0-Flash (team2023gemini) after adding a randomly selected atomic visual capability. The distribution of k in the dataset shifts to the right by 1, effectively increasing the average complexity of the training dataset. Fig. 1 shows that VIT on higher k dataset leads to higher performance on downstream tasks, after controlling for the data generation method. These results motivate our complexity-aware VIT data curation method.
Figure 1: Complexity k. We show that increasing the complexity of LLaVA-665K improves performance. (left) Examples of questions with different k-values, where k is the number of atomic capabilities required. (middle) Distribution of k-value in VIT subset (LLaVA) and VIT subset augmented with 1 additional capability (LLaVAk+1). (right) Performance on downstream tasks (§4.1) for VIT subset (LLaVA), VIT subset regenerated with no capability augmentation (LLAVAk+0) and VIT subset augmented with 1 additional capability (LLaVAk+1).
We propose COMPACT (COMPositional Atomic-to-complex Visual Capability Tuning), an efficient VIT data recipe that controls the complexity of each sample by combining atomic visual capabilities. We summarize our key contributions:
  1. 1.
    We show that increasing the complexity of training samples allow more effective use of information content in a given dataset. We introduce COMPACT to address limitations in complexity-agnostic scaling of conventional VIT methods.
  2. 2.
    We define the $k$-value to quantify the complexity of a vision-language task. We analyze the optimal distribution of complexity in the data for maximum efficiency and performance.
  3. 3.
    We demonstrate the effectiveness of COMPACT. With only 10% volume of the LLaVA-665K (llava15) VIT dataset, training with our COMPACT data matches the performance of full-scale VIT (100.2% relative performance), and even outperforms on particularly complex multimodal benchmarks like MM-Vet (mmvet) (29.2 when trained with full LLaVA-665K vs 31.7 when trained with COMPACT) and MMStar  (mmstar) (35.1 vs. 36.1).
@@
Table 1: Taxonomy of atomic capabilities. We identify 10 atomic capabilities that are necessary for visual reasoning. We categorize them into attribution, recognition, and relation. Atomic capabilities serve as building blocks for building complex tasks.
@@
2 Related Work
Visual Instruction Tuning.  Instruction following is an essential capability in language models (wei2021finetuned; zhou2023instruction). Misalignment between a model’s response and the format requested by a question can hinder the precise evaluation of its performance (he2024does; hsieh2023sugarcrepe; salido2025none; balepur2025these). VIT involves training a model on a fixed set of response patterns that can be repeated during inference (e.g., multiple-choice, short- and long-response questions) (liu2023visual; llava15). Although VIT has shown performance improvements in general multimodal capabilities (huang2023visual), recent work (ghosh2024closer) has shown that optimizing for response formatting potentially limits the quality of language model responses.
While VIT (llava15) relies on simple questions to train instruction following capability, COMPACT leverages more complex questions to achieve the same goal. Higher $k$ questions facilitate learning by encouraging the model to learn from more visual features in each image.
Complex Tasks in LLMs and MLLMs.  Complexity in the space of vision-language tasks has been loosely conceptualized by different interpretations of compositionality. Some studies view compositionality as a sequential arrangement of basic tasks (chen2024facilitating; li2024mosaic). Compositionality as integrations of basic capabilities has been examined in geometric reasoning  (chae24decomposing) and general visual reasoning (hua2024mmcomposition), but mainly in the context of evaluation.
We argue that compositional complexity can be studied in the context of training MLLMs. Studies highlight that while MLLMs do show signs of compositional capability  (ossowski2024prompting), they struggle when constituting components and their combined patterns are not strongly learned or missing during training  (campbell2025understanding). COMPACT builds on this insight by defining complexity as the combination of atomic visual capabilities and widening the range of complexity represented in the dataset.
Scaling Visual Instruction Tuning Data.  VIT is a data- and compute-heavy step in training (xu2023parameter). Prior work has approached this challenge from mainly two directions. First, upscaling-based approaches (cvbench; llava15) argue that ever larger corpora can continue to improve visual capabilities, partially addressing the challenge of low information density through sheer volume of data. Second, data selection studies argue that full performance can be reproduced with smaller amounts of data (lee2024concept; liu2024less). ICONS (wu2024icons) shows that models can achieve near-full performance across a suite of MLLM benchmarks using only a fraction of the original VIT dataset.
However, these approaches treat complexity as a byproduct of scale rather than a controllable metric. COMPACT exposes the model to more complex tasks without scaling the dataset size by explicitly targeting higher $k$-values during sample generation.
3 Method
We aim to construct a training dataset in which each sample requires knowledge of the fundamental visual capabilities and their meaningful combinations that progressively increase the complexity of the task. Inspired by the compositional nature of visual reasoning, we first define complexity as a metric that arises from combining basic visual capabilities. This requires identification of a set of fundamental visual capabilities that can be combined, called atomic visual capabilities (§3.1). A subset of these capabilities are combined to generate a new sample with target complexity, which corresponds to COMPACT’s four-step data recipe (§3.2).
3.1 Atomic Visual Capabilities
Atomic capabilities are foundational skills that can be combined to solve complex tasks. For example, a model needs to acquire object recognition, color attribution, and spatial relationship understanding capabilities to identify how objects of different colors are spatially oriented. For each task $T$, we identify a set of atomic visual capabilities $\{c_{1},\dots c_{k}\}$ required to solve this task. We define the number of atomic capabilities required to solve the task $T$ as its complexity $k$.
Figure 2: COMPACT data generation pipeline. (Left): We design a data recipe that can scale the complexity of each training example. We randomly sample $k_{gen}\in\{1,2,3\}$ atomic capabilities such as color, object recognition, and spatial relationship. (Center): We generate questions that integrate all $k_{gen}$ sampled capabilities and verify their quality. (Right): We combine the synthetically generated compositional tuning data with instruction tuning data for response formatting.
We build the taxonomy of atomic capabilities (Tab. 1) based on two key considerations. First, we focus on vision-centric skills that can encourage the model utilize the visual information in the image. Therefore, we exclude non-perceptual capabilities (e.g. cultural knowledge, historical context, coding, and math) that require external information beyond the image. Second, we include capabilities that can also be found in existing VIT data curation methods (liu2023visual; cvbench; li2025eagle), for fair comparison of our approach to others. As a result, we create a list of 10 atomic capabilities which can be divided into three major categories: 1) Attribution: identifying visual properties (e.g., color and shape). 2) Recognition: identifying visual entities, including objects, actions, and text. 3) Relation: identifying visual interactions and spatial orientations.
We note that the atomic capabilities in Tab. 1 are not expected to cover the entire multimodal task space or be completely orthogonal (see Appendix Fig. 8). Rather, we aim to find sufficiently distinct capabilities that allow meaningful combinations to generate diverse visual reasoning tasks. We provide more details and discussion on the taxonomy curation in Appendix §A.4.
@@ 3.2 Visual Compositional Tuning Data Recipe
COMPACT generates multi-capability questions $\mathcal{D}_{\text{comp}}$ by prompting vision-language models to create questions that require natural [^1] We use the term “natural” to denote combination of visual capabilities that correspond to their co-occurrence patterns in real-world settings, wherein multiple capabilities are integrated in a way that is contextually and semantically meaningful. integration of multiple atomic visual capabilities. This process involves four key steps: (1) sampling a subset of capabilities from our predefined set of atomic visual capabilities (Capability Sampling), (2) prompting Gemini-2.0-Flash (team2023gemini) to generate questions that integrate all the selected capabilities (Conversation Generation), (3) validating the capability requirement and the quality of each question (Quality Verification), and (4) combining our generated compositional tuning data with a small portion of the LLaVA-665K (llava15) VIT data for response formatting (Dataset Assembly).
@@
Step 1: Capability Sampling.  We start by taking a random sample of images from LLaVA-665K (llava15). For each image, we repeatedly sample $k_{gen}\in\{1,2,3\}$ capabilities from our predefined pool of 10 atomic visual capabilities. At this stage, $k_{gen}$ functionally serves as the lower bound of the actual $k$-value of the generated conversation. We sample $k_{gen}$ with the expectation that the final $k$-value of the sample will eventually be higher ($k_{gen}\leq k$), as some atomic capabilities are weakly correlated in practice (see Fig. 8 in Appendix). In order to maximize capability coverage, we do the following in each round of capability sampling: (a) prioritize the capabilities that have not been selected for that image, and (b) drop duplicate combinations of capabilities for the same image. These efforts ensure that our training examples capture diverse visual information.
@@
Step 2: Conversation Generation.  For each capability combination that is sampled, we prompt Gemini-2.0-Flash (team2023gemini) to generate a conversational question-answer pair that integrates all capabilities in the combination, as well as a score between 0 and 100 that represents its confidence in the quality of the conversation. Our carefully designed prompt (see Appendix §B) enforces several key constraints: (a) questions must require the use of visual information from the image and cannot be answered from its text alone, (b) answers must be concise, (c) questions must integrate the specified capabilities naturally (without using conjunctions to simply conjoin single-capability questions), and (d) questions must reference objects and features actually present in the image. The purpose of these constraints is to produce vision-centric conversations that are unambiguous and natural.
@@
Step 3: Quality Verification.  We include a verification process with Gemini-2.0-Flash (team2023gemini) to ensure the quality and diversity of the training dataset. We filter out questions with uninformative answers (e.g., “unknown”, “not visible”) or those with confidence scores below 70%. We discard questions that share more than 60% of the words with those previously accepted.
We then perform capability verification by prompting Gemini-2.0-Flash (team2023gemini) to analyze whether each question indeed requires the $k_{gen}$ specified capabilities. The generation and verification processes in steps 2 and 3 repeat iteratively until we collect 2-3 high-quality conversations per $k_{gen}$ for each image or reach a maximum of 10 verification attempts.
Step 4: Dataset Assembly.  We address the challenge of aligning the response format by mixing our synthetically generated compositional tuning data with a random 5% subset of the LLaVA-665K (llava15) VIT dataset. This mixture of instruction tuning and compositional tuning data has the following effects. First, the VIT subset maintains the model’s ability to handle diverse response formats and instructions required by modern MLLM benchmarks (e.g., multiple-choice questions (mme), open-ended answers (llava15)). Second, our compositional tuning data leverages complex questions to facilitate learning with higher information density. In this way, we delegate the instruction following capability training to the original VIT data and allow our compositional tuning data to focus on visual reasoning.
The size of the compositional tuning data is determined by the minimum number of images needed to match full LLaVA-665K (llava15) VIT performance. COMPACT preserves the contents of the images, which enables us to fairly compare against existing methods in their ability to extract rich visual information.
Table 2: Baseline comparisons. COMPACT outperforms baseline VIT data recipes on multimodal benchmarks. With only 5% of the LLaVA-665K (llava15) VIT data and 32K of our compositional tuning data (65K total), COMPACT outperforms the random subset of the VIT data (Random), various data reduction methods, as well as the full VIT data. The best and second best results for each benchmark are shown in bold and underlined, respectively. We further provide results using an open-source generator (Qwen3-VL-4B-Instruct) for COMPACT in Table 9.
@@
Recipe # Data InfoVQA SeedBench2Plus MME TextVQA $M$-Vet CV-Bench MMStar LLaVA-W Rel. (%)
LLaVA-665K 665K 20.80 41.72 1478.48 46.99 29.22 60.92 35.11 68.50 100.00
@@
EL2N 65K 20.52 42.95 1378.58 42.41 33.53 50.92 33.82 66.40 97.09
Perplexity 65K 20.46 41.90 1375.32 42.95 30.32 52.72 33.47 68.40 96.09
SemDeDup 65K 20.54 43.96 1431.36 42.71 28.39 42.24 34.18 62.10 93.31
D2-Pruning 65K 20.90 43.70 1343.34 41.82 31.61 48.49 36.63 68.40 97.13
Self-Sup 65K 20.61 42.51 1434.30 42.68 30.18 54.30 34.33 61.20 96.04
Self-Filter 65K 19.99 41.33 1290.34 36.59 28.21 45.17 33.67 66.40 90.51
ICONS 65K 21.0 42.03 1402.75 43.12 31.23 55.96 35.96 61.8 97.47
@@
Figure 3: Performance across compositional tuning data scales. We show that COMPACT’s compositional tuning data scales more efficiently than conventional VIT. We fix the VIT subset (5% of LLaVA-665K (llava15)) and scale the compositional tuning data in COMPACT from 2K to 32K. We compare each mix with VIT only datasets with equal data budgets. COMPACT (solid lines) consistently outperforms LLaVA-665K VIT (dashed lines) with fewer data. COMPACT’s compositional tuning data scales particularly well on SeedBench2Plus (seedbench), which consists of spatially complex tasks of navigating charts and maps.